OpenStudy (anonymous):

Solve the following inequality and write your answer using interval notation. Please show all of your work. 50/x+4 +1<4x

OpenStudy (anonymous):

You need to get an equivalent inequality with 0 on one side \[\frac{50}{x+4}+1<4x\] \[\frac{50}{x+4}+1-4x<0\] \[\frac{50+(x+4)-4x(x+4)}{x+4}<0\] \[\frac{-4x^2-15x+54}{x+4}<0\]

OpenStudy (anonymous):

Now you have a critical value of x=-4 You must also solve the zeros of the numerator. This will divide the x-axis into intervals. All values in each interval will be either positive or negative You're looking for the intervals which result in a negative

OpenStudy (anonymous):